OpenStudy (anonymous):

find a formula for the inverse of the function 32 Fahrenheit =0C f(x)9/5x+32Fahrenheit

ganeshie8 (ganeshie8):

cuold you format the question correctly ? if psble please take a screenshot of original question and attach

ganeshie8 (ganeshie8):

\(\large f(x) = \dfrac{9}{5}x + 32\)

ganeshie8 (ganeshie8):

basically this formula converts temperature in Celcius to Fahrenheit

ganeshie8 (ganeshie8):

you need to find the inverse to convert from Fahrenheit to Celcius

ganeshie8 (ganeshie8):

do you know how to find the inverse of a function ?

OpenStudy (anonymous):

k -32 5/9x=y

ganeshie8 (ganeshie8):

very good try!

ganeshie8 (ganeshie8):

\(\large f(x) = \dfrac{9}{5}x + 32\) subtract 32 both sides : \(\large f(x)-32 = \dfrac{9}{5}x \) multiply both sides by 5/9 : \(\large \dfrac{5}{9}(f(x)-32) = x \) replace x and f(x) : \(\large \dfrac{5}{9}(x-32) = f(x) \)

ganeshie8 (ganeshie8):

So the inverse funciton for converting Fahrenheit to Celcius is \(\large f(x) = \dfrac{5}{9}(x-32) \)

ganeshie8 (ganeshie8):

plugin x=32 above to to convert 32 fahrenheit to celicius

OpenStudy (anonymous):

so i plug in 32 for x and get 5/9

ganeshie8 (ganeshie8):

do you really get 5/9 when you plugin x=32 ?

OpenStudy (anonymous):

o never mind 0

OpenStudy (anonymous):

multiply everything by 5/9

ganeshie8 (ganeshie8):

\(\large f(x) = \dfrac{5}{9}(x-32)\) plugin x=32 : \(\large f(32) = \dfrac{5}{9}(32-32) = \dfrac{5}{9}(0) = 0\)

ganeshie8 (ganeshie8):

so the given statement is is indeed True : 32 Fahrenheit = 0 C

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

yw!